OpenStudy (anonymous):

If 1^3x-5/ 16 = 1 / 2^3-x , Find x.

OpenStudy (anonymous):

Help?

OpenStudy (mertsj):

\[\frac{1^{3x-5}}{16}=\frac{1}{2^{3-x}}\]

OpenStudy (mertsj):

Is that the question?

OpenStudy (anonymous):

find x.

OpenStudy (mertsj):

If so, we know that 1 raised to any power is 1 so the numerators are both 1. Therefore, the denominators must be equal so: \[2^4=2^{3-x}\]

OpenStudy (mertsj):

And 4 = 3-x so x = -1

OpenStudy (anonymous):

Great! Thanks! :)

OpenStudy (mertsj):

yw